# SemanticFormer: Holistic and Semantic Traffic Scene Representation for Trajectory Prediction using Knowledge Graphs

### v2

@@ Abstract
Trajectory prediction in autonomous driving relies on accurate representation of all relevant contexts of the driving scene, including traffic participants, road topology, traffic signs, as well as their semantic relations to each other. Despite increased attention to this issue, most approaches in trajectory prediction do not consider all of these factors sufficiently. We present SemanticFormer, an approach for predicting multimodal trajectories by reasoning over a semantic traffic scene graph using a hybrid approach. It utilizes high-level information in the form of meta-paths, i.e. trajectories on which an agent is allowed to drive from a knowledge graph which is then processed by a novel pipeline based on multiple attention mechanisms to predict accurate trajectories. SemanticFormer comprises a hierarchical heterogeneous graph encoder to capture spatio-temporal and relational information across agents as well as between agents and road elements. Further, it includes a predictor to fuse different encodings and decode trajectories with probabilities. Finally, a refinement module assesses permitted meta-paths of trajectories and speed profiles to obtain final predicted trajectories. Evaluation of the nuScenes benchmark demonstrates improved performance compared to several SOTA methods. In addition, we demonstrate that our knowledge graph can be easily added to two graph-based existing SOTA methods, namely VectorNet and Laformer, replacing their original homogeneous graphs. The evaluation results suggest that by adding our knowledge graph the performance of the original methods is enhanced by 5% and 4%, respectively.
@@ I INTRODUCTION
Autonomous vehicles are recognized as a promising solution to address critical challenges such as road safety, traffic congestion, and energy optimization. A crucial task towards the realization of autonomous driving vision is motion prediction. It involves determining a set of spatial coordinates that represent the predicted movement of a given agent within a future time window. However, motion prediction is a challenging task due to various contextual factors such as the difficulty of intention prediction, the complex interactions of traffic participants, the intricate road topology, comprising lanes, lane dividers, and pedestrian crossings, as well as adherence to traffic regulations. State-of-the-art approaches utilize various representations for traffic scenes such as raster-based [1, 2], or graph-based [3, 4] to capture and utilize contextual information sufficiently.
@@
  • A hybrid architecture with attention mechanisms to model the semantic relationships and dependencies between traffic agents and road elements for accurate multi-modal trajectory prediction.
  • Extensive evaluation of our method on the nuScenes dataset [6] including several ablation studies.
@@ II RELATED WORK
Methods that use heterogeneous graphs, i.e. graphs with different entity types such as vehicles, bicycles or pedestrians and relation types like agent-to-lane or agent-to-agent, are recently proposed [17, 18, 19, 20, 21, 22]. However, they are limited to only a portion of the relevant information and are unable to fully capture all scene details and the interconnections between the entities. Our approach aims to fill this gap using formal ontologies for constructing a knowledge graph [23, 24] capturing the rich information of traffic scenes. We also define different domain-specific relation types between nodes with relation-specific features to represent the relational information between nodes.
@@ IV-E4 Effect of Heterogeneous Graph Operators
We analyze different heterogeneous graph operators like HGT [59] and HAN [55]. As shown in Table VII, to prevent overfitting, we merge sub-classes like single solid, double solid, etc, to switchViaPermitted and switchViaNonPermitted relationships to represent lane-changing situations.

### v3

@@ Abstract
Trajectory prediction in autonomous driving relies on accurate representation of all relevant contexts of the driving scene, including traffic participants, road topology, traffic signs, as well as their semantic relations to each other. Despite increased attention to this issue, most approaches in trajectory prediction do not consider all of these factors sufficiently. We present SemanticFormer, an approach for predicting multimodal trajectories by reasoning over a semantic traffic scene graph using a hybrid approach. It utilizes high-level information in the form of meta-paths, i.e. trajectories on which an agent is allowed to drive from a knowledge graph which is then processed by a novel pipeline based on multiple attention mechanisms to predict accurate trajectories. SemanticFormer comprises a hierarchical heterogeneous graph encoder to capture spatio-temporal and relational information across agents as well as between agents and road elements. Further, it includes a predictor to fuse different encodings and decode trajectories with probabilities. Finally, a refinement module assesses permitted meta-paths of trajectories and speed profiles to obtain final predicted trajectories. Evaluation of the nuScenes benchmark demonstrates improved performance compared to several SOTA methods. In addition, we demonstrate that our knowledge graph can be easily added to two graph-based existing SOTA methods, namely VectorNet and LaFormer, replacing their original homogeneous graphs. The evaluation results suggest that by adding our knowledge graph the performance of the original methods is enhanced by 5% and 4%, respectively. Graph data is available at https://github.com/boschresearch/nuScenes_Knowledge_Graph
@@ I INTRODUCTION
Autonomous vehicles are recognized as a promising solution to address critical challenges such as road safety, traffic congestion, and energy optimization. A crucial task towards the realization of autonomous driving vision is motion prediction[1]. It involves determining a set of spatial coordinates that represent the predicted movement of a given agent within a future time window. However, motion prediction is a challenging task due to various contextual factors such as the difficulty of intention prediction, the complex interactions of traffic participants, the intricate road topology, comprising lanes, lane dividers, and pedestrian crossings, as well as adherence to traffic regulations. State-of-the-art approaches utilize various representations for traffic scenes such as raster-based [2, 3], or graph-based [4, 5] to capture and utilize contextual information sufficiently.
@@
  • A hybrid architecture with attention mechanisms to model the semantic relationships and dependencies between traffic agents and road elements for accurate multi-modal trajectory prediction. Evaluated on nuScenes dataset [7].
@@ II RELATED WORK
Methods that use heterogeneous graphs, i.e. graphs with different entity types such as vehicles, bicycles or pedestrians and relation types like agent-to-lane or agent-to-agent, are recently proposed [18, 19, 20, 21, 22, 23]. However, they are limited to only a portion of the relevant information and are unable to fully capture all scene details and the interconnections between the entities. Our approach aims to fill this gap using formal ontologies for constructing a knowledge graph [24, 25, 26, 27, 28] capturing the rich information of traffic scenes. Knowledge graphs have been applied in other automotive applications like POI recommendation [29, 30] and driving situation understanding [31].
@@ IV-E4 Effect of Heterogeneous Graph Operators
We analyze different heterogeneous graph operators like HGT [67] and HAN [63]. As shown in Table VII, to prevent overfitting, we merge sub-classes like single solid, double solid, etc, to switchViaPermitted and switchViaNonPermitted relationships. *N means number of layers of operator is N.